# Seeding with Differentially Private Network Information

### v2

@@ 2.1 Influence sampling and maximization
In the independent cascade model, infections spread via an edge-weighted undirected graph ${\mathcal{G}}=(V,p)$ on $n$ nodes. The set of nodes, $V$, is indexed by $[n]:=\{1,\ldots,n\}$ and $p:V\times V\to[0,1]$ assigns probability weights to every edge in ${\mathcal{G}}$. An infection begins at a set $S$ of seed nodes and spreads through the graph in rounds. Each infected node, once they become infected, has a single chance of passing the infection to their neighbors. Each undirected edge $e=(v,u)$ has a weight $p_{e}\in[0,1]$ representing the probability that the process spreads along edge $e$ to node $u$ in the round immediately after $v$ is first infected.
The above process has the following equivalent formulation: We first sample an undirected, unweighted graph $g$ by independently realizing each edge $e$ with probability $p_{e}$. We write $g\sim{\mathcal{G}}$ to denote that $g$ is drawn from the random graph distribution ${\mathcal{G}}$. Then we can construct the set of infected nodes in the independent cascade process, as the set of nodes that are reachable from seed set $S$ in $g$. Now we define some notations. Given an undirected graph $g$ and subsets of nodes $S$ and $T$, we write $C_{g}(S)$ for the set of nodes that are reachable from (or infected by) $S$ in $g$, and $C_{g}(S|T):=C_{g}(S)\setminus C_{g}(T)$ for the set of nodes that are reachable from $S$ but not from $T$. Then the influence function $I_{g}(S):=|C_{g}(S)|$ is the number of nodes reachable (or infected) by $S$ and $I_{\mathcal{G}}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$ is its expectation. Similarly, we define the conditional influence function $I_{g}(S|T):=|C_{g}(S|T)|$ and denote its expected value by $I_{\mathcal{G}}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S|T)]$.
@@
How can we design a seeding algorithm $M$ that takes $m$ influence samples ${\mathbf{x}}$ and outputs a set of $k$ nodes, $M({\mathbf{x}})\in\mathcal{Y}:=\{0,1\}^{V}$? Here we present a seeding algorithm by Eckles et al. (2022). The core idea of Algorithm 1 is to estimate the influence function from the influence samples, and run the greedy algorithm. Specifically, given $m$ realized influence samples ${\mathbf{x}}$ and a subset of nodes $S$ in ${\mathcal{G}}$, we define the $S$-reachable influence samples as: $C_{\mathbf{x}}(S):=\{t\in[m]:\exists v\in S,x_{v}^{t}=1\},$ which is the set of influence samples whose activated nodes are reachable from the set $S$, and $C_{\mathbf{x}}(S|T)=C_{\mathbf{x}}(S)\setminus C_{\mathbf{x}}(T)$. Alternatively, if we treat ${\mathbf{x}}$ a collection of $m$ subset $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$. Finally, we define $I_{\mathbf{x}}(S)=\frac{n}{m}|C_{\mathbf{x}}(S)|$ and $I_{\mathbf{x}}(S|T)=\frac{n}{m}|C_{\mathbf{x}}(S|T)|$. The following proposition says the expectation of $I_{\mathbf{x}}$ is equal to $I_{{\mathcal{G}}}$. We will use $v$ as $\{v\}$ for singleton sets to simplify the notation, e.g., $I_{\mathcal{G}}(v)$ as $I_{\mathcal{G}}(\{v\})$.
@@
Given a positive integer $m$, a graph ${\mathcal{G}}=(V,p)$ and sets of nodes $S,T\subseteq V$, if ${\mathbf{x}}$ is the random variable of $m$ influence sample, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{\mathcal{G}}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{\mathcal{G}}(S|T)$. Additionally, by Hoeffding’s inequality, for all $\epsilon>0$ we have: $\Pr[|I_{{\mathbf{x}}}(S)-I_{\mathcal{G}}(S)|\geq\epsilon n]\leq 2\exp(-2\epsilon^{2}m),$ where the probability is over the randomness in the generative model of influence samples $x$.
With the above observation, Eckles et al. (2022) design a seeding algorithm that only uses influence samples and show that if the number of influence sample is of order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to $\alpha n$ additive error. Informally, Algorithm 1 is a greedy algorithm that sequentially query Algorithm 2 for the seed that maximizes the marginal improvement. In this paper, we will follow similar idea that combine greedy algorithm and an estimation algorithm for marginal improvement.
@@
In Algorithm 1, we slightly modify the INF-SAMPLE$(m,k)$ algorithm to sequentially call Algorithm 2 for selecting the next seed. Instead of collecting fresh influence samples at each greedy step, Algorithm 1 calls Algorithm 2 to reuse the influence samples and have much better influence sample complexity as a result. We formally show this improvement in LABEL:thm:exp:acc.
@@
Given ${\mathcal{G}}$ of $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
It is well known that even with the full knowledge of the graph ${\mathcal{G}}$ no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP. Eckles et al. (2022) show that without the knowledge of ${\mathcal{G}}$, Algorithm 1 can approach an optimal efficient algorithm when the number of influence samples is large enough. In this paper, we ask if we can achieve similar performance when privacy constraints are imposed. Eckles et al. (2022) also show that one cannot hope to provide a constant factor approximation guarantee using $o(n)$ influence samples.
@@
Let $10<\alpha<$ 1 be any constant. For any algorithm with $o(n)$ samples, there exists a weighted graph ${\mathcal{G}}$ so that the output of the algorithm can only infect less than $\alpha\mbox{\sf OPT}$ nodes.
@@ 2.2 Influence Sample Differential Privacy
Intuitively, a differentially private (DP) algorithm ensures that the algorithm outputs on two adjacent inputs are close to each other, making them hard to distinguish. However, there are multiple alternative definitions of adjacency relation on graph-structured input data. One classical notion is node DP where two graphs are adjacent if they are different by one node’s connections. Another one is edge DP where two graphs are adjacent if they differ by one edge (Hay et al. 2009). In this paper, we would like to focus on DP seeding algorithms that use influence samples but not raw graph data. Instead of raw graph, we study algorithms that only use historical cascades data — i.e., influence samples — where two collections of $m$ influences samples ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$ in $\{0,1\}^{n\times m}$ are (influence sample) adjacent if we can change one entry to convert $\mathbf{x}$ to $\mathbf{x}^{\prime}$, and write ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$ if they are adjacent. Given the above adjacency relation on influence samples, we can define if a stochastic seeding algorithm $M$ is DP.
@@
Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ at one row (i.e., an individual can decide whether to report their whole history or not). Our analysis in the paper still holds after strengthening the privacy constraints from $\epsilon$ to $\epsilon/m$; please refer to Section 6 for a more detailed discussion.
@@
We believe the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016). First, the notion of node DP allows an agent to misreport her connection in the social networks. However, when collecting data on network cascades, a single agent can hardly modify her connection to change the dynamics of social contagion, but rather she can decide whether to misreport her previous contagion data. Therefore, we argue that defining adjacent datasets based on influence samples as the input, rather than the raw social network data, is more natural and realistic. Moreover, we show in LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg that node-DP and edge-DP are too strict for influence maximization and hardly admit any utility guarantees. In Section 4 we will elaborate on the nature of our privacy protections in social network data collection and intervention design contexts.
@@
Give two edge-weighted undirected graphs ${\mathcal{G}}=(V,p)$ and ${\mathcal{G}}^{\prime}=(V^{\prime},p^{\prime})$, we say ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exists $v\in V$, $p_{e}=p_{e}^{\prime}$ for all $v\notin e$. Similarly, ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$. For $k<n$, we say that a seeding algorithm $M^{k}$ is $\epsilon$-node (or edge) DP if for any $k$, node (or edge) adjacent graphs ${\mathcal{G}},{\mathcal{G}}^{\prime}$, and output ${\mathbf{y}}$, $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$.
@@
The following proposition (proved in Section 7.1) shows no node DP seeding algorithm can output a high quality seed.
@@
Proposition 2.5 (label = prop:node:dp:neg, name=Hardness of Influence Maximization with Node-DP)
@@
Given $n\geq 1$ and $\epsilon>0$, for any $\epsilon$-node-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists ${\mathcal{G}}$ of $n$ nodes so that for $k=1$ we have: $\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1).$
@@
We use the fact that a star and a totally isolated graph are different by one node’s connection. A node-DP algorithm needs to output similar seed sets on these two graphs, and thus returns low quality seeds. Based on similar intuition, the following proposition (proved in Section 7.2) shows no edge-DP seeding algorithm can output a high quality seed either. Both results can be strengthened to exclude the possibility of useful algorithms with “approximate” $(\epsilon,\delta)$-DP guarantees as well. The latter is a relaxation of the defining constraint in Equation 1 to allow for an additive term $\delta$; see Section 7.2.
@@
Proposition 2.6 (label = prop:edge:dp:neg, name=Hardness of Influence Maximization with Edge-DP)
@@
For all $n\geq 1$ and $\epsilon>0$, there exists $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$ so that any $\epsilon$-edge-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists graph ${\mathcal{G}}$ on $n$ nodes so that for $k=1$ we have: $(1-1/e)\textstyle\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
@@
Compared to LABEL:thm:original, the above results show that no node-DP (or edge-DP) algorithm can perform close to $(1-1/e)\mbox{\sf OPT}$ even with the full knowledge of ${\mathcal{G}}$. In Section 7.2 (Proposition 7.1), we show that similar lower bounds hold even if we consider an additive relaxation of the node- and edge-DP requirements, allowing the $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor. In this paper, we will design influence sample DP algorithms that perform comparable to Algorithm 1. Finally, note that node-DP implies edge-DP, but influence sample DP is not comparable to node- or edge-DP because the algorithms take different inputs.
@@
Our algorithms will make use of the following two general DP mechanisms (we use algorithm and mechanism interchangeably): exponential mechanism (McSherry and Talwar 2007, Dwork et al. 2014), and randomized respond mechanism (Warner 1965, Evfimievski et al. 2003). First, to achieve $\epsilon$-DP, given a finite set (range), $\mathcal{R}$, the exponential mechanism is defined with respect to a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ and its sensitivity $\Delta(u):=\max_{r\in\mathcal{R},{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},r)-u({\mathbf{x}}^{\prime},r)|$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ as defined in Dwork et al. (2014). Moreover, as shown in the following theorem, for a fixed dataset ${\mathbf{x}}$ exponential mechanism’s output almost maximizes the scoring function $u({\mathbf{x}},r)$ and can be interpreted as a DP soft max.
@@
Consider a dataset ${\mathbf{x}}\in\{0,1\}^{n\times m}$, privacy budget $\epsilon>0$, a finite set $\mathcal{R}$, and a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ with sensitivity $\Delta(u)$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ that outputs an element $r\in\mathcal{R}$ with probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ is $\epsilon$-ISDP and for all $t>0$, we have:
@@
The above exponential mechanism ensures that the output does not reveal too much information about the privacy-critical input data, limiting how much an adversary can infer about the input data according to Definition 2.4. However, depending on data curation requirements (who can access the exact user data), protection against output-based inferences may not be sufficient and we may need to limit the algorithm’s access to exact input data. Such privacy guarantees can be formulated as local DP where each individual’s data is privatized before running the algorithms. One standard locally DP algorithm is the randomized response mechanism (Evfimievski et al. 2003). Specifically, an $\epsilon$-DP randomized response mechanism guarantees privacy by independently flipping each bit of the input dataset (influence sample collection) with probability $1/(e^{\epsilon}+1)$.

### v3

@@ 2.1 Influence sampling and maximization
In the independent cascade model, infections spread via an edge-weighted undirected graph ${\mathcal{G}}=(V,p)$ on $n$ nodes. The set of nodes, $V$, is indexed by $[n]:=\{1,\ldots,n\}$ and $p:V\times V\to[0,1]$ assigns probability weights to every edge in ${\mathcal{G}}$. An infection begins in a set $S$ of seed nodes and spreads through the graph in rounds. Each infected node, once it becomes infected, has a single chance of passing the infection to their neighbors. Each undirected edge $e=(v,u)$ has a weight $p_{e}\in[0,1]$ that represents the probability that the process spreads along edge $e$ to node $u$ in the round immediately after $v$ is first infected.
Some notation.
The above process has the following equivalent formulation: we first sample an undirected, unweighted graph $g$ by independently realizing each edge $e$ with probability $p_{e}$. We write $g\sim{\mathcal{G}}$ to denote that $g$ is drawn from the random graph distribution ${\mathcal{G}}$. Then we can construct the set of infected nodes in the independent cascade process, as the set of nodes that can be reached from the seed set $S$ in $g$. Given an undirected graph $g$ and subsets of nodes $S$ and $T$, we write $C_{g}(S)$ for the set of nodes that are reachable from (or infected by) $S$ in $g$, and $C_{g}(S|T):=C_{g}(S)\setminus C_{g}(T)$ for the set of nodes that are reachable from $S$ but not from $T$. The influence function $I_{g}(S):=|C_{g}(S)|$ is the number of nodes reachable (or infected) by $S$ and $I_{\mathcal{G}}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$ is its expectation. Similarly, we define conditional influence function $I_{g}(S|T):=|C_{g}(S|T)|$ and denote its expected value by $I_{\mathcal{G}}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S|T)]$.
@@
How can we design a seeding algorithm $M$ that takes $m$ influence samples ${\mathbf{x}}$ and produces a set of $k$ nodes, $M({\mathbf{x}})\in\mathcal{Y}:=\{0,1\}^{V}$? Here we present a seeding algorithm by Eckles et al. (2022). The core idea of Algorithm 1 is to estimate the influence function from the influence samples and to run the greedy algorithm. Specifically, given $m$ realized influence samples ${\mathbf{x}}$ and a subset of nodes $S$ in ${\mathcal{G}}$, we define the $S$-reachable influence samples as: $C_{\mathbf{x}}(S):=\{t\in[m]:\exists v\in S,x_{v}^{t}=1\},$ which is the set of influence samples whose activated nodes are reachable from the set $S$, and $C_{\mathbf{x}}(S|T)=C_{\mathbf{x}}(S)\setminus C_{\mathbf{x}}(T)$. Alternatively, if we treat ${\mathbf{x}}$ as a collection of $m$ subsets $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$. Finally, we define $I_{\mathbf{x}}(S)=\frac{n}{m}|C_{\mathbf{x}}(S)|$ and $I_{\mathbf{x}}(S|T)=\frac{n}{m}|C_{\mathbf{x}}(S|T)|$. The following proposition states that the expectation of $I_{\mathbf{x}}$ is equal to $I_{{\mathcal{G}}}$. We will use $v$ instead of $\{v\}$ for singleton sets to simplify the notation, for example, writing $I_{\mathcal{G}}(v)$ for $I_{\mathcal{G}}(\{v\})$.
@@
Given a positive integer $m$, a graph ${\mathcal{G}}=(V,p)$, and sets of nodes $S,T\subseteq V$, if ${\mathbf{x}}$ is the random variable that encodes $m$ influence samples, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{\mathcal{G}}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{\mathcal{G}}(S|T)$. Additionally, by Hoeffding’s inequality, for all $\epsilon>0$ we have: $\Pr[|I_{{\mathbf{x}}}(S)-I_{\mathcal{G}}(S)|\geq\epsilon n]\leq 2\exp(-2\epsilon^{2}m),$ where the probability is over the randomness in the generative model of influence samples $x$.
With the above observation, Eckles et al. (2022) design a seeding algorithm that only uses influence samples and show that if the number of influence samples is of the order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to the additive error $\alpha n$. Informally, Algorithm 1 is a greedy algorithm that sequentially queries Algorithm 2 for the seed that maximizes marginal improvement. In this paper, we will follow a similar idea that combines a greedy algorithm and an estimation algorithm for marginal improvement.
@@
In Algorithm 1, we slightly modify the INF-SAMPLE$(m,k)$ algorithm to sequentially call Algorithm 2 to select the next seed. Instead of collecting fresh influence samples at each greedy step, Algorithm 1 calls Algorithm 2 to reuse the influence samples and, as a result, to have a much better influence sample complexity. We formally show this improvement in LABEL:thm:exp:acc.
@@
Given the graph ${\mathcal{G}}$ on $n\geq 2$ nodes, we want to seed $k\leq n$ nodes. For any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
It is well known that even with complete knowledge of the graph ${\mathcal{G}}$, no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP. Eckles et al. (2022) show that without the knowledge of ${\mathcal{G}}$, Algorithm 1 can approach an optimal efficient algorithm when the number of influence samples is large enough. In this paper, we ask if we can achieve similar performance when privacy constraints are imposed. Eckles et al. (2022) also show that one cannot hope to provide a constant factor approximation guarantee using $o(n)$ influence samples.
@@
Let $10<\alpha<$ 1 be any constant. For any algorithm with $o(n)$ samples, there exists a weighted graph ${\mathcal{G}}$ so that the output of the algorithm can only infect fewer than $\alpha\mbox{\sf OPT}$ nodes.
@@ 2.2 Influence Sample Differential Privacy
Intuitively, a differentially private (DP) algorithm ensures that the algorithms output on two adjacent inputs are close to each other, making them difficult to distinguish. However, there are multiple alternative definitions of the adjacency relation in graph-structured input data. One classical notion is node DP, where two graphs are adjacent if they are different by the connections of one node. Another is edge DP, where two graphs are adjacent if they differ by one edge (Hay et al. 2009). In this paper, we focus on developing differentially private (DP) seeding algorithms that utilize influence samples, rather than relying on raw graph data, which is often impractical to obtain in sensitive contexts such as public health interventions for HIV prevention. In such cases, contact tracing using surveys of sexual activity may be preferable to collecting a complete network due to privacy concerns. The highly contextualized and variable nature of contact data makes a reliable network representation not well defined, even if it is desirable. Therefore, we abstract away form the network representation and work directly with influence samples, which, in the case of contact tracing for HIV prevention, track the spread of sexual activity over time without requiring complete network information.
We study algorithms that only use these influence samples, where two collections of $m$ influence samples, ${\mathbf{x}}$ and ${\mathbf{x}}^{\prime}$ in $\{0,1\}^{n\times m}$, are “influence sample” adjacent if one entry can be modified to transform ${\mathbf{x}}$ into ${\mathbf{x}}^{\prime}$. We denote this adjacency by ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$. Given this adjacency relation on influence samples, we define a stochastic seeding algorithm $M$ to be DP if the algorithm’s output does not change significantly when applied to adjacent influence samples. In the HIV example, by focusing on influence samples derived from sexual behavior, our DP algorithms provide privacy guarantees for participants in sensitive public health settings while identifying high-risk individuals for targeted interventions.
@@
Instead of entry-level adjacency, one may consider individual-level adjacency that two inputs are adjacent if they differ in one row (that is, an individual can decide whether to report their entire history or not). Our analysis in the paper still holds after strengthening the privacy constraints from $\epsilon$ to $\epsilon/m$; please refer to Section 6 for a more detailed discussion.
@@
We believe that the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (Kasiviswanathan et al. 2013, Raskhodnikova and Smith 2016). First, the notion of node DP allows an agent to misreport her social network connections. However, when collecting data on network cascades, a single agent can hardly modify her connections to change the contagion dynamics but instead can decide whether to misreport her contagion data. Therefore, we argue that defining adjacent datasets based on influence samples as input, rather than raw social network data, is more natural and realistic for contact tracing and public health applications. Moreover, we show in LABEL:prop:node:dp:neg and LABEL:prop:edge:dp:neg that node DP and edge DP are too strict for influence maximization and hardly admit any utility guarantees. In Section 4, we will elaborate on the nature of our privacy protections in social network contexts, including for data collection and intervention design.
@@
Given two edge-weighted undirected graphs ${\mathcal{G}}=(V,p)$ and ${\mathcal{G}}^{\prime}=(V^{\prime},p^{\prime})$, we say ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exist $v\in V$, such that $p_{e}=p_{e}^{\prime}$ for all $e\notniFromTxfonts v$. Similarly, ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ such that $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$. For $k<n$, we say that a seeding algorithm $M^{k}$ is $\epsilon$-node (or edge) DP if for any $k$, node (or edge) adjacent graphs ${\mathcal{G}},{\mathcal{G}}^{\prime}$, and output ${\mathbf{y}}$, $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$.
@@
The following proposition (proved in Section 7.1) shows that seeding algorithms with node DP guarantees cannot produce high-quality seeds.
@@
Proposition 2.5 (label = prop:node:dp:neg, name=Hardness of Influence Maximization with Node DP)
@@
Given $n\geq 1$ and $\epsilon>0$, for any $\epsilon$ node-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists ${\mathcal{G}}$ of $n$ nodes so that for $k=1$ we have: $\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1).$
@@
We use the fact that a star and a totally isolated graph are different by one node’s connection. A node-DP algorithm needs to output similar seed sets on these two graphs and thus returns low-quality seeds. Based on a similar intuition, the following proposition (proved in Section 7.2) shows that no edge-DP seeding algorithm can output a high-quality seed either. Both results can be strengthened to exclude the possibility of useful algorithms with “approximate” $(\epsilon,\delta)$-DP guarantees as well. The latter is a relaxation of the defining constraint in Equation 1 to allow for an additive term $\delta$; see Section 7.2.
@@
Proposition 2.6 (label = prop:edge:dp:neg, name=Hardness of Influence Maximization with Edge DP)
@@
For all $n\geq 1$, $\epsilon>0$, and any $\epsilon$ edge-DP seeding algorithm $M^{k}({\mathcal{G}})$, there exists a graph ${\mathcal{G}}$ on $n$ nodes so that for $k=1$ and some $\alpha_{\epsilon}\geq\frac{1}{1000e^{\epsilon}}$ we have: $(1-1/e)\textstyle\max_{S:|S|=1}I_{\mathcal{G}}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}({\mathcal{G}})}[I_{\mathcal{G}}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
@@
Compared to LABEL:thm:original, the above results show that node-DP (or edge-DP) algorithms cannot perform close to $(1-1/e)\mbox{\sf OPT}$ even with full knowledge of ${\mathcal{G}}$. In Section 7.2 (Proposition 7.1), we show that similar lower bounds hold even if we consider additive relaxations of the node and edge DP, allowing the $\Pr[M^{k}({\mathcal{G}})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}({\mathcal{G}}^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor. In this paper, we will design influence sample DP algorithms that perform comparable to Algorithm 1. Finally, note that node DP implies edge-D , but influence sample DP is not comparable to node or edge DP because the algorithms take different inputs.
@@
Our algorithms make use of the following two general DP mechanisms (we use algorithm and mechanism interchangeably): exponential mechanism (McSherry and Talwar 2007, Dwork et al. 2014), and randomized respond mechanism (Warner 1965, Evfimievski et al. 2003). First, to achieve $\epsilon$-DP, given a finite set (range), $\mathcal{R}$, exponential mechanism is defined with respect to a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ and its sensitivity $\Delta(u):=\max_{r\in\mathcal{R},{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},r)-u({\mathbf{x}}^{\prime},r)|$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ produces an element $r\in\mathcal{R}$ with a probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ as defined in Dwork et al. (2014).
Moreover, as shown in the following theorem, for a fixed data set ${\mathbf{x}}$, the output of the exponential mechanism almost maximizes the scoring function $u({\mathbf{x}},r)$ and can be interpreted as a softmax that satisfies the DP randomization requirement.
@@
Consider a data set ${\mathbf{x}}\in\{0,1\}^{n\times m}$, privacy budget $\epsilon>0$, a finite set $\mathcal{R}$, and a scoring function $u:\{0,1\}^{n\times m}\times\mathcal{R}\to{\mathbb{R}}$ with sensitivity $\Delta(u)$. The exponential mechanism $\mathcal{M}^{\epsilon}_{E}({\mathbf{x}},u,\mathcal{R})$ that produces an element $r\in\mathcal{R}$ with a probability proportional to $\exp\left(\frac{\epsilon u({\mathbf{x}},r)}{2\Delta(u)}\right)$ is $\epsilon$-ISDP and for all $t>0$, we have:
@@
The above exponential mechanism ensures that the output does not reveal too much information about the privacy-critical input data, limiting how much an adversary can infer about the input data according to Definition 2.4. However, depending on data curation requirements (who can access the exact user data), protection against output-based inferences may not be sufficient, and we may need to limit the algorithm’s access to exact input data. Such privacy guarantees can be formulated as local DP where the data of each individual are privatized before running the algorithms. A standard locally DP algorithm is randomized response mechanism (Evfimievski et al. 2003): an $\epsilon$-DP randomized response mechanism guarantees privacy by independently flipping each bit of the input data set (influence sample collection) with probability $1/(e^{\epsilon}+1)$.